OpenStudy (anonymous):

What is the square root of 72x^3 divided by y^2

OpenStudy (anonymous):

\[\sqrt{72x ^3 \div y^2}\]

OpenStudy (anonymous):

\[\frac{ 432\sqrt{2} }{ y }\]

OpenStudy (anonymous):

72=9*8 =3*3*2*2*2 =(3^2)*(2^2)*2 now [\sqrt{72}\]=3*2*[\sqrt{2}\]

OpenStudy (anonymous):

@judispute what is this ??

OpenStudy (anonymous):

so exactly what is the anser of this question khurramshahzad 72=9*8 =3*3*2*2*2 =(3^2)*(2^2)*2 now [\sqrt{72}\]=3*2*[\sqrt{2}\]

OpenStudy (anonymous):

(6/y)*[\sqrt{2x}\]

OpenStudy (anonymous):

sqrt(x3)=x*sqrt(x) sqrt(y2)=y

OpenStudy (anonymous):

(6/y)*[\sqrt{2x}\] its answer